OpenStudy (anonymous):

The expression (cscx + cotx)^2 is the same as _____. A. csc^2x + cot^2x B. 1 + 2cot^2x + 2cscx cotx C. 1 + 2csc^2x D. csc^2x + 2cscx + cot^2x

OpenStudy (anonymous):

U might as well just feed these things straight into Wolfram...

OpenStudy (dumbcow):

\[(\frac{1}{\sin x} + \frac{\cos x}{\sin x})^{2} = (\frac{1+\cos x}{\sin x})^{2} = \frac{\cos^{2} x + 2\cos x + 1}{\sin^{2} x} = \cot^{2} x +2\cot xcsc x +\csc^{2} x\]

OpenStudy (dumbcow):

its close to D but not exact, are the answer options correct

myininaya (myininaya):

cow you did it the long way lol

OpenStudy (dumbcow):

yeah im wierd like that

OpenStudy (anonymous):

so id go with d?

myininaya (myininaya):

don't go with any of them none of them are correct unless you mistyped something

OpenStudy (anonymous):

everythings right

myininaya (myininaya):

then write none of these and put what cow got

myininaya (myininaya):

wait

myininaya (myininaya):

\[1+\cot^2x=\csc^2x\] \[\cot^2x+2cotxcscx+\csc^2x=\cot^2x+2cotxscscx+1+\cot^2x=2\cot^2x+2cotxcscx+1\]

OpenStudy (anonymous):

Problem is u can often write these things in a number of ways, takes some time to solve...

myininaya (myininaya):

i was wrong at first

OpenStudy (dumbcow):

wait its B)...yeah i just got that too

myininaya (myininaya):

i'm terrible

myininaya (myininaya):

we could had used that identity 1+cot^2x=csc^2x

OpenStudy (dumbcow):

trig can be tricky cause you can write the same thing different ways